Solve the system. Enter the solution coordinates.
2x - 4y = 7
4x + 2y = 9

Answer by funmath(2933)

I am going to show you how to solve this system using what some books call the elimination method and others the addition method. If you are supposed to use some other method, resubmit your question stating so.
We want to be able to add these two equations together and eliminate one of the variables. This can be done a couple of ways. I'm going to multiply the top equation by -2.
2x-4y=7
4x+2y=9
----------
-2(2x-4y=7)
4x+2y=9
----------
-4x+8y=-14
4x+2y=9
-----------
Add the two equations together:
-4x+8y=-14
+4x+2y=9
___________
0+10y=-5
solve for y:
10y=-5
10y/10=-5/10
y=-1/2
Substitute -1/2 into the original top equation for y and solve for x.
2x-4(-1/2)=7
2x+2=7
2x+2-2=7-2
2x=5
2x/2=5/2
x=5/2
The coordinates (x,y)=(5/2,-1/2)